OpenStudy (anonymous):

\(\color{green}{\textsf {Integrate :}}\) \[\large \color{blue}{\int\limits \frac{x^2}{(x^2 + (z-c)^2)^\frac{3}{2}} dx}\]

hartnn (hartnn):

z and c are constants ?

OpenStudy (anonymous):

What substitution or rearrangement I should make here??? Yes, z and c are constants...

OpenStudy (kainui):

Looks like a trig substitution, pretend (z-c) is just a single thing.

OpenStudy (anonymous):

That is surely not a problem, we can take (z-c) as some constant k also..

OpenStudy (kainui):

I agree @waterineyes but I don't know what's giving you problems or not. The key to knowing what to choose for your trig substitution is just to look to see if anything is of this form: \[\Large \sin^2 \theta +\cos^2 \theta = 1\] See we can rearrange this to go from two things to just a single square to remove square roots simpler: \[\Large \cos^2 \theta = 1-\sin^2 \theta \]

hartnn (hartnn):

adjust the numerator x^2 = (x^2 +k^2) - k^2 then separate into 2 terms

hartnn (hartnn):

see if that approach is shorter than any substitution....?

hartnn (hartnn):

you can use the direct formula for 1/sqrt (x^2+a^2) right ?

OpenStudy (anonymous):

I try using that rearrangement....

OpenStudy (anonymous):

For one term I can use, but I think next term is again creating problem for me..

OpenStudy (anonymous):

I want to ask if the integral would be : \[\int\limits \frac{dx}{(x^2 + k^2)^{\frac{3}{2}}}\]

OpenStudy (kainui):

I have a feeling that will be rough. Let's just divide the pythagorean identity by cos^2 theta to get: \[\Large \frac{\sin^2 \theta}{\cos^2 \theta}+\frac{\cos^2 \theta}{\cos^2 \theta}=\frac{1}{\cos^2 \theta}\] this gives us: \[\Large \tan^2 \theta + 1 = \sec^2 \theta\] which can be rearranged to give: \[\Large \tan^2 \theta = \sec^2 \theta - 1\] Ok so now combine these last two with the original pythagorean identity and we essentially have a 1 to 1 correspondence between what we should substitute. \[\Large y^2=1-x^2 \\ \Large y^2= 1+x^2 \\ \Large y^2 =-1+x^2\]

OpenStudy (anonymous):

I think the next integral would be that I have shown above.. Yes??? @hartnn

OpenStudy (kainui):

So for example if we have: \[\Large \int\limits \frac{dx}{\sqrt{a^2-x^2}}dx\] \[\Large x=a \sin \theta\]

hartnn (hartnn):

correct! \(\int\limits \frac{dx}{(x^2 + k^2)^{\frac{3}{2}}}\) a u-substitute could work here....if you feel this is getting complicated, go for trig

OpenStudy (anonymous):

@Kainui, try to solve asked integral using trig substitution.. I am unable to think of any substitution here..

OpenStudy (anonymous):

@hartnn yes integral is getting more complex and calculating if I do that, but I try.. Please you think of any other simpler one to reach there..

OpenStudy (kainui):

I have solved it, but you haven't really responded to me trying to help you out. Read what I wrote and I'll help you understand it, but I'm not going to give you the answer. \[\Large \int\limits \frac{x^2}{(x^2+(z-c)^2)^{3/2}}dx\]

OpenStudy (anonymous):

@Kainui Hey don't give me the answer, actually I have one.. But I want to know how to reach there...

hartnn (hartnn):

\(\int\limits \frac{dx}{(x^2 + k^2)^{\frac{3}{2}}}\) is itself easy, but yes, we finally have to do atrig substitution for this i think x = a tan y will work

OpenStudy (anonymous):

The example you are giving, there trig substitution is easy to be applied.. I want to know any subs that will work with asked integral..

hartnn (hartnn):

x= k tan y i mean

OpenStudy (anonymous):

dx = ksec^2(y).dy

hartnn (hartnn):

correct, denominatro, you will have sec^3

OpenStudy (anonymous):

\[\frac{k \sec^2(y).dy}{\sec^3(y)} = \frac{\cos(y)}{k^2}\]

OpenStudy (anonymous):

Yep, in second integral we can use substitution which is working successfully... :)

OpenStudy (anonymous):

Now I solve it in neat and clean way... :) Thanks to you @hartnn and @Kainui

hartnn (hartnn):

welcome ^_^